Suppose the sum of the 151st term and the 251st term of an arithmetic sequence is 10. Find the sum of the first 401 terms of the sequence. Provide 3 examples of such a sequence.
Answer by venugopalramana(3286)

1. Suppose the sum of the 151st term and the 251st term of an
arithmetic sequence is 10. Find the sum of the first 401 terms of the
sequence. Provide 3 examples of such a sequence.
T151=A+(151-1)D=A+150D....................I
T251=A+(251-1)D=A+250D....................II
EQN.II+EQN.I..GIVES
2A+400D=10........................................III
S401=(401/2){2A+(401-1)D}=401*10/2=2005
EX.I.....PUT A=205...D=-1...SO....
205,204,203,202,.......................
EX.II.......PUT A=405...D=-2....SO
405,403,401......
EX.III......PUT A=605....D=-3.....SO
605,602,599.........